OpenStudy (anonymous):

sin(x+2pi/3)+sin(x-2pi/3)=-rad2/2 Solve for 0 lessthanorequalto x less than 2pi

OpenStudy (anonymous):

\[\sin(x +2pi/3 ) + \sin(x - 2pi/3) = \sqrt{2}/2\]

OpenStudy (anonymous):

\[Solve for 0 \le x < 2pi\]

OpenStudy (anonymous):

hint, rewrite the above as \(-\sin(x)\) and then solve

OpenStudy (anonymous):

\[= - \sqrt{2}/2\]

OpenStudy (anonymous):

thats what it equals i wrote it wrong

OpenStudy (anonymous):

did you solve for x?

OpenStudy (anonymous):

no i mean i wrote the original equation wrong and didnt place a negative in front

OpenStudy (anonymous):

ok so you have \[-\sin(x)=-\frac{\sqrt{2}}{2}\] \[\sin(x)=\frac{\sqrt{2}}{2}\] solve for x

OpenStudy (anonymous):

two solutions would be \(x=\frac{\pi}{4}\) and \(x=\frac{3\pi}{4}\)

OpenStudy (anonymous):

noo i dont have that.... i have - sinx + sin 2pi/3 -sinx + sin 2pi/3 = rad2/2

OpenStudy (anonymous):

what do i do from there?

OpenStudy (anonymous):

I have 2sinx = -rad2/2 what do i do now?

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

welli have to solve for in radians

OpenStudy (anonymous):

0 <or equal x < 2pi

hero (hero):

satellite should come back here and post a more complete explanation. I keep telling reminding him not to half-arse his explanations.

OpenStudy (anonymous):

im not friends with him, so i cant make him come back

hero (hero):

He's offline and most likely done for the night.

OpenStudy (anonymous):

are you paid to do this?

hero (hero):

I get paid the same amount as you do.

OpenStudy (anonymous):

i dont get apid for this ohhh so you dont haha

hero (hero):

I can get the proper solution to this for you

hero (hero):

I am useful in that regard.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

bye

hero (hero):

Where are you going?